# LLM-Based Multi-Hop Question Answering with Knowledge Graph Integration in Evolving Environments

### v1

@@ Abstract
The rapid obsolescence of information in Large Language Models (LLMs) has driven the development of various techniques to incorporate new facts. However, existing methods for knowledge editing still face difficulties with multi-hop questions that require accurate fact identification and sequential logical reasoning, particularly among numerous fact updates. To tackle these challenges, this paper introduces Graph Memory-based Editing for Large Language Models (GMeLLo), a straitforward and effective method that merges the explicit knowledge representation of Knowledge Graphs (KGs) with the linguistic flexibility of LLMs. Beyond merely leveraging LLMs for question answering, GMeLLo employs these models to convert free-form language into structured queries and fact triples, facilitating seamless interaction with KGs for rapid updates and precise multi-hop reasoning. Our results show that GMeLLo significantly surpasses current state-of-the-art knowledge editing methods in the multi-hop question answering benchmark, MQuAKE, especially in scenarios with extensive knowledge edits.
@@ 1 Introduction
As the widespread deployment of LLMs continues, the imperative to keep their knowledge accurate and up-to-date, without incurring extensive retraining costs, becomes increasingly evident Sinitsin et al. (2020). Several approaches have been proposed in prior works to address this challenge, with some focusing on the incremental injection of new facts into language models Rawat et al. (2020); De Cao et al. (2021); Meng et al. (2022); Mitchell et al. (2022a). Interestingly, certain methodologies in the literature diverge from the conventional path of updating model weights, opting instead for an innovative strategy involving the use of external memory to store the edits Mitchell et al. (2022b); Zhong et al. (2023).
As LLMs operate as black boxes, modifying one fact might inadvertently alter another, making it challenging to guarantee accurate revisions. In this paper, we introduce GMeLLo, an effective approach designed to synergize the strengths of LLMs and KGs in addressing the multi-hop question answering task after knowledge editing Zhong et al. (2023). An illustrative example of our focus is presented in Figure  1. Following an update regarding the information of the British Prime Minister, it becomes evident that the corresponding spouse information should also be modified.
As depicted in Figure  2, our GMeLLo comprises the following key steps:
@@
LLMs, trained on extensive sentence corpora Brown et al. (2020); Rae et al. (2022); Chowdhery et al. (2023), are expected to encapsulate a wide range of commonly used sentence structures. As a result, they are invaluable tools for analyzing sentences and extracting entities and relations. Once the correct relation chain and edited triples are obtained, using a formal query to interrogate the KG in a Knowledge-based Question Answering (KBQA) Cui et al. (2017) manner ensures precision in the searching process. In cases where KBQA fails, we still have LLMs for question answering (QA) to ensure comprehensive coverage. GMeLLo outperforms current state-of-the-art (SOTA) methods on two datasets from the MQuAKE benchmark, affirming its effectiveness in multi-hop question answering within an evolving environment.
@@ 2.2 Knowledge Editing
2.2.1 Modifying Model Parameters
In the case of modifying model parameters, this can be further categorized into meta-learning or locate-and-edit approaches. Meta-learning methods, as discussed in De Cao et al. (2021); Mitchell et al. (2022a), utilize a hyper network to learn the necessary adjustments for editing LLMs. The locate-then-edit paradigm, as demonstrated in Dai et al. (2022); Meng et al. (2022, 2023); Li et al. (2023a); Gupta et al. (2023); Zhang et al. (2024), involves initially identifying parameters corresponding to specific knowledge and subsequently modifying them through direct updates to the target parameters.
2.2.2 Preserving Model Parameters
In the case of preserving model parameters, the introduction of additional parameters or external memory becomes necessary. The paradigm of additional parameters, as presented in Dong et al. (2022); Hartvigsen et al. (2022); Huang et al. (2022), incorporates extra trainable parameters into the language model. These parameters are trained on a modified knowledge dataset, while the original model parameters remain static. On the other hand, memory-based models Mitchell et al. (2022b); Zhong et al. (2023) explicitly store all edited examples in memory and employ a retriever to extract the relevant edit facts for each new input, guiding the model in generating the edited output.
@@ 3.1 Extracting Fact Triples from Edited Information Using LLMs
In our approach, we leverage Wikidata Vrandečić and Krötzsch (2014), a widely recognized KG, as the foundational knowledge base. When updated facts are received, we utilize LLMs to extract entities from the sentences and determine their relationships (selecting a relation from the predefined list). This process generates edited fact triples, which are then used to update the KG (see Figure  2). Updating the KG with an edited fact triple involves identifying the connections in the KG based on the subject entity and relation, breaking these connections, and establishing a new connection based on the triple.
@@ 3.2 Extracting Relation Chain from Questions Using LLMs
As the world evolves rapidly, the training data for LLMs can quickly become outdated. However, since the evolution of linguistic patterns typically progresses at a slower pace, the extensive training data of LLMs should enable them to effectively comprehend most sentence patterns. In this paper, we employ LLMs to extract the relation chain from a sentence, encompassing the mentioned entity in the question and its relations with other unidentified entities. Similar to the fact triple exaction 3.1, we task LLMs with selecting a relation from a predefined list to mitigate varied representations of the same relation. Take a question sentence from the MQuAKE-CF Zhong et al. (2023) dataset as an example,
@@
The presented question necessitates a 4-hop reasoning process. With "Eeyore" as the known entity in focus, the journey to the final answer involves identifying its creator ’?x’, moving on to the creator’s child ’?y’, obtaining the child’s country of citizenship ’?z’, and culminating with the retrieval of the country’s capital ’?m’. All the relations, such as ’creator,’ ’child,’ ’country of citizenship,’ and ’capital,’ are chosen from a predefined list of relations. The relation chain encapsulates all essential information for deriving the answer.
@@
3.3 Converting Relation Chain into a Formal Query
Once the relation chain is obtained, the next step involves integrating the known entity and the relations into a formal query template. Consider a KG represented in RDF [^2] https://www.w3.org/RDF/ format and a corresponding SPARQL [^3] https://www.w3.org/TR/sparql11-query/ query, the relation chain elucidated in Section 3.2 should be represented as follows,
@@ 3.3 Converting Relation Chain into a Formal Query
In this context, "ent" and "rel" serve as prefixes for entity and relation, respectively. The identifier "E0" uniquely represents "Eeyore" within the KG, while the identifiers for "creator," "child," "country of citizenship," and "capital" are denoted as "R0", "R1", "R2", and "R3" respectively. After identifying the entity "?id", we retrieve its string label "?label" as the final answer.
@@ 3.4 Integrating LLM-based QA and KBQA
However, addressing multi-hop questions, especially those where the edited facts pertain to intermediary hops, presents a challenge in accurately retrieving the relevant information and performing correct multi-hop question answering. This challenge is particularly pronounced when dealing with a large volume of edited facts. For instance, accurately identifying the relevant fact given the question in Figure  2 and producing the correct final answer is difficult.
KBQA. To address the challenges of LLM-based question answering, we integrate responses from KBQA to refine the outputs from the LLMs, as detailed in Sections 3.1-3.3. When the relation chain and fact triples are accurately derived, the KBQA system provides the correct answer. However, if the relation chain is incorrectly extracted, the search path in the KG may become invalid, leading the KBQA system to yield no output. In such instances, we accept the response from the LLMs as the final answer.
@@ 4 Experiment
In the upcoming section, we will conduct experiments to demonstrate the effectiveness of employing our GMeLLo methodology.
@@ 4.1.4 Knowledge Graph Setting
Considering Wikidata’s community-driven nature, guaranteeing a dynamic and comprehensive dataset across a spectrum of knowledge domains, we use Wikidata Vrandečić and Krötzsch (2014) as the foundational KG for this experiment. To align the relations in the question and fact sentences with those in WikiData Vrandečić and Krötzsch (2014), we follow the following steps:
  • We select the first 500 item properties [^5] https://www.wikidata.org/w/index.php?title=Special:ListProperties/wikibase-item&limit=500&offset=0 from WikiData as the base relations. Items represent either concrete or abstract entities, such as a person Piscopo and Simperl (2019).
  • Next, we employ GPT-3.5-Turbo to examine each multi-hop question in the test samples to determine if it contains any of the base relations.
  • Afterward, we rank the frequencies of each relation and choose the top 50 relations as candidates for use in relation chain extraction and edited fact triple extraction.
@@ 4.2 Main Results
As shown in Table 1, our GMeLLo significantly outperforms all existing methods on the both the MQuAKE-CF dataset and the MQuAKE-T dataset Zhong et al. (2023), particularly when handling a large number of edits.
In addition, we evaluated MeLLo and GMeLLo using two larger models, GPT-3.5-Turbo-Instruct and GPT-3.5-Turbo [^8] https://platform.openai.com/docs/models/gpt-3-5-turbo, on the MQuAKE-CF dataset with k=3000999The model text-davinci-003 used in Zhong et al. (2023) was deprecated on January 4, 2024.. The accuracy rates achieved by MeLLo and GMeLLo with GPT-3.5-Turbo-Instruct were 30.7% and 51.4%, respectively. While GMeLLo achieved an accuracy of 66.4% with GPT-3.5-Turbo, the same model consistently returned errors when tested with MeLLo, suggesting that the prompts may require modification for compatibility with chat completion models. These results indicate that GMeLLo performs well even when scaled to larger LLMs.
@@ 4.3 Ablation Study
As demonstrated in Table  2, the performance of KBQA remains consistent, as all edited facts are converted to triples and all relation chains are extracted from the test questions, regardless of the value of ’k’. However, as the parameter ’k’ increases, more edited facts are stored in the external memory. Consequently, selecting the relevant edits to accurately answering the questions becomes increasingly challenging for LLM-based QA.
@@ 4.3.1 Further Analysis
To evaluate the impact of KBQA on LLM-based QA within the GMeLLo framework, we conducted an analysis comparing the responses from LLMs to those from the KG. We consider the KG’s response as the final answer. Therefore, comparing to only using LLM-based QA, if the answer from LLMs is correct but the answer from the KG is incorrect, this leads to a decline in performance. Conversely, if the answer from LLMs is incorrect but the answer from the KG is correct, performance improves. If the KBQA provides no response, performance remains unchanged. As illustrated in Table 3, when there are discrepancies between KBQA and LLM-based QA responses, the likelihood of KBQA providing the correct answer increases as the parameter k increases.
@@
GPT-J-6B ${\color[rgb]{1,0,0}\boldsymbol{\uparrow}}{}$ 8.1 22.9 24.9 25.0 44.0 47.2 47.9 48.0
${\color[rgb]{0,1,0}\boldsymbol{\downarrow}}{}$ 12.5 2.4 1.2 0.7 0.7 0.4 0.3 0.3
@@
Vicuna-7B ${\color[rgb]{1,0,0}\boldsymbol{\uparrow}}{}$ 7.7 17.8 19.6 20.0 4.2 19.7 21.4 21.7
${\color[rgb]{0,1,0}\boldsymbol{\downarrow}}{}$ 21.8 3.9 2.0 1.2 7.2 4.2 4.0 3.9
@@ 4.4 Qualitative Analysis
Table  2 illustrates that Vicuna exhibits superior performance in directly handling the QA task, particularly when provided with the exact edited facts. Conversely, GPT-J excels in sentence analysis tasks, showcasing its high performance in the KBQA task.
@@ 4.4.1 Inferior Performance of GPT-J in QA
Table 2 shows that the performance of GPT-J and Vicuna in conducting QA tasks is comparable on the MQuAKE-CF dataset when k=1. However, GPT-J exhibits notably lower performance on the MQuAKE-T dataset. Further analysis revealed that GPT-J struggles in answering questions with only an edited fact pertaining to its intermediary information, such as: {examplebox} Sample from MQuAKE-CF
@@ 5 Conclusion
In this paper, we present GMeLLo, a method designed for multi-hop question answering in dynamic environments. Except leveraging LLMs for question answering, we also leverage the capabilities of LLMs to extract the triples from edited fact sentence to update KG, and use the capabilities of LLMs to analyze question sentences and generate a relation chain, and finally get the formal query by filling in a formal query template. Finally, we combine KBQA and LLM-based QA to bolster the multi-hop question answering capability within a dynamic environment. This approach capitalizes on the strengths of both LLMs and KGs. By utilizing LLMs for analyzing question sentences and QA to ensure the coverage, and KBQA to provide accurate results, we achieve a synergy between these two methodologies.
@@ A.2 Hyperparmers Settings
To ensure reproducibility, we set the temperature to zero in all experiments. Table 5 shows that retrieving the top-6 edited facts from external memory provides the best average performance on the MQuAKE-CF dataset for $k>1$. Consequently, we include top-6 edited facts in the prompt for subsequent experiments on this dataset when $k>1$. Similarly, for the MQuAKE-T dataset when $k>1$, we opted to incorporate the top-1 edited fact in the prompt.

### v2

@@ Abstract
The important challenge of keeping knowledge in Large Language Models (LLMs) up-to-date has led to the development of various methods for incorporating new facts. However, existing methods for such knowledge editing still face difficulties with multi-hop questions that require accurate fact identification and sequential logical reasoning, particularly among numerous fact updates. To tackle these challenges, this paper introduces Graph Memory-based Editing for Large Language Models (GMeLLo), a straightforward and effective method that merges the explicit knowledge representation of Knowledge Graphs (KGs) with the linguistic flexibility of LLMs. Beyond merely leveraging LLMs for question answering, GMeLLo employs these models to convert free-form language into structured queries and fact triples, facilitating seamless interaction with KGs for rapid updates and precise multi-hop reasoning. Our results show that GMeLLo significantly surpasses current state-of-the-art (SOTA) knowledge editing methods in the multi-hop question answering benchmark, MQuAKE, especially in scenarios with extensive knowledge edits.
@@ 1 Introduction
An important challenge in deploying Large Language Models (LLMs) is keeping their knowledge accurate and up-to-date, without incurring expensive retraining costs Sinitsin et al. (2020). Several approaches have been proposed in prior works to address this challenge. Some methods focus on the incremental injection of new facts into language models Rawat et al. (2020); De Cao et al. (2021); Meng et al. (2022); Mitchell et al. (2022a). Alternatively, other methods involve the use of external memory to store new facts Mitchell et al. (2022b); Zhong et al. (2023), which does not require updating LLM model weights.
As LLMs operate as black boxes, modifying one fact might inadvertently alter another, making it challenging to guarantee accurate revisions. In this paper, we introduce GMeLLo, an effective approach designed to synergize the strengths of LLMs and Knowledge Graphs (KGs) in addressing the multi-hop question answering task after knowledge editing Zhong et al. (2023). An illustrative example is presented in Figure 1. Following an information update regarding the British Prime Minister, it becomes evident that the corresponding spouse information should also be modified.
As depicted in Figure 2, our GMeLLo method comprises the following key steps:
@@
LLMs, trained on extensive sentence corpora Brown et al. (2020); Rae et al. (2022); Chowdhery et al. (2023), are expected to encapsulate a wide range of commonly used sentence structures. As a result, they are invaluable tools for analyzing sentences and extracting entities and relations. Once the correct relation chain and edited triples are obtained, using a formal query to interrogate the KG in a Knowledge-based Question Answering (KBQA) Cui et al. (2017) manner ensures precision in the searching process. In cases where KBQA fails, we still have LLMs for question answering (QA) to ensure comprehensive coverage. GMeLLo outperforms current SOTA methods on two datasets from the MQuAKE benchmark, affirming its effectiveness in multi-hop question answering within an evolving environment.
@@ 2.2 Knowledge Editing
2.2.1 Parameter-Modification Paradigm
In the case of modifying model parameters, this can be further categorized into meta-learning or locate-and-edit approaches. Meta-learning methods De Cao et al. (2021); Mitchell et al. (2022a) utilize a hyper network to learn the necessary adjustments for editing LLMs. The locate-then-edit paradigm Dai et al. (2022); Meng et al. (2022, 2023); Li et al. (2023a); Gupta et al. (2023); Zhang et al. (2024) involves initially identifying parameters corresponding to specific knowledge and subsequently modifying them through direct updates to the target parameters.
2.2.2 Parameter-Preservation Paradigm
In the case of preserving model parameters, the introduction of additional parameters or external memory becomes necessary. The paradigm of additional parameters Dong et al. (2022); Hartvigsen et al. (2022); Huang et al. (2022) incorporates extra trainable parameters into the language model. These parameters are trained on a modified knowledge dataset, while the original model parameters remain static. In contrast, memory-based models Mitchell et al. (2022b); Zhong et al. (2023); Gu et al. (2024) explicitly store all edited examples in memory and employ a retriever to extract the relevant edit facts for each new input, guiding the model in generating the updated output.
@@ 3.1 Extracting Fact Triples from Edited Information Using LLMs
In our approach, we leverage Wikidata Vrandečić and Krötzsch (2014), a widely recognized KG, as the foundational knowledge base. When updated facts are received, we utilize LLMs to extract entities from the sentences and determine their relationships (selecting a relation from the predefined list). This process generates edited fact triples, which are then used to update the KG (see Figure 2). Updating the KG with an edited fact triple involves identifying the connections in the KG based on the subject entity and relation, breaking these connections, and establishing a new connection based on the triple.
@@ 3.2 Extracting Relation Chain from Questions Using LLMs
As the world evolves rapidly, the training data for LLMs can quickly become outdated. However, since the evolution of linguistic patterns typically progresses at a slower pace, the extensive training data of LLMs should enable them to effectively comprehend most sentence patterns. In this paper, we employ LLMs to extract the relation chain from a sentence, encompassing the mentioned entity in the question and its relations with other unidentified entities. Similar to the fact triple exaction mentioned in Section 3.1, we task LLMs with selecting a relation from a predefined list to mitigate varied representations of the same relation. Take a question sentence from the MQuAKE-CF Zhong et al. (2023) dataset as an example,
@@
The presented question necessitates a 4-hop reasoning process. With "Eeyore" as the known entity in focus, the journey to the final answer involves identifying its creator "?x", moving on to the creator’s child "?y", obtaining the child’s country of citizenship "?z", and culminating with the retrieval of the country’s capital "?m". All the relations, such as "creator", "child", "country of citizenship", and "capital", are chosen from a predefined list of relations. The relation chain encapsulates all essential information for deriving the answer.
@@
3.3 Converting a Relation Chain into a Formal Query
Once the relation chain is obtained, the next step involves integrating the known entity and the relations into a formal query template. For a KG represented in RDF [^2] https://www.w3.org/RDF/ format, the relation chain elucidated in Section 3.2 can be represented as the following SPARQL [^3] https://www.w3.org/TR/sparql11-query/ query,
@@ 3.3 Converting a Relation Chain into a Formal Query
In this context, "ent" and "rel" serve as prefixes for entity and relation, respectively. The identifier "E0" uniquely represents "Eeyore" within the KG, while the identifiers for "creator," "child," "country of citizenship," and "capital" are denoted as "R0", "R1", "R2", and "R3", respectively. After identifying the entity "?id", we retrieve its string label "?label" as the final answer.
@@ 3.4 Integrating LLM-based QA and KBQA
However, addressing multi-hop questions, especially those where the edited facts pertain to intermediary hops, presents a challenge in accurately retrieving the relevant information and performing correct multi-hop question answering. This challenge is particularly pronounced when dealing with a large volume of edited facts. For instance, accurately identifying the relevant fact given the question in Figure 2 and producing the correct final answer is difficult.
KBQA. To address the challenges of LLM-based question answering, we integrate responses from KBQA to refine the outputs from the LLMs, as detailed in the previous section. When the relation chain and fact triples are accurately derived, the KBQA system provides the correct answer. However, if the relation chain is incorrectly extracted, the search path in the KG may become invalid, leading the KBQA system to yield no output. In such instances, we accept the response from the LLMs as the final answer.
@@ 4 Experiment
In this section, we will present the results from our experiments to demonstrate the effectiveness of employing our GMeLLo methodology.
@@ 4.1.3 Baselines
  • PokeMQA Gu et al. (2024). It also uses a memory-based approach, which decouples question decomposition from knowledge editing to reduce the burden on LLMs. Additionally, it introduces auxiliary knowledge prompts to assist with question decomposition.
@@ 4.1.4 Knowledge Graph Setting
Considering Wikidata’s community-driven nature, guaranteeing a dynamic and comprehensive dataset across a spectrum of knowledge domains, we use Wikidata Vrandečić and Krötzsch (2014) as the foundational KG for this experiment. To align the relations in the question and fact sentences with those in WikiData Vrandečić and Krötzsch (2014), we take the following steps:
@@ 4.2 Main Results
As shown in Table 1999We use the baseline performance reported in Zhong et al. (2023) and Gu et al. (2024). Since the experiment settings in Gu et al. (2024) differ from those in Zhong et al. (2023), we only include the results from Gu et al. (2024) under the same settings. A dash (’-’) indicates that performance was not reported for that setting., our GMeLLo significantly outperforms all existing methods on the both the MQuAKE-CF dataset and the MQuAKE-T dataset Zhong et al. (2023), particularly when handling a large number of edits.
@@
To further validate our findings with more capable base models, we evaluated MeLLo and GMeLLo using two larger models, GPT-3.5-Turbo-Instruct and GPT-3.5-Turbo, on the MQuAKE-CF dataset with k=3000101010The model text-davinci-003 used in Zhong et al. (2023) was deprecated on January 4, 2024.. The accuracy rates achieved by MeLLo and GMeLLo with GPT-3.5-Turbo-Instruct were 30.7% and 51.4%, respectively. While GMeLLo achieved an accuracy of 66.4% with GPT-3.5-Turbo, the same model consistently returned errors when tested with MeLLo, suggesting that the prompts may require modification for compatibility with chat completion models. These results indicate that GMeLLo performs well even when scaled to larger LLMs.
@@ 4.3 Ablation Study
As demonstrated in Table 2, the performance of KBQA remains consistent because all edited facts are converted to triples and all relation chains are extracted from the test questions, regardless of the value of "k". Correctly answering a multi-hop question in KBQA requires both accurate extraction of fact triples and the relation chain. However, as the parameter "k" increases, more edited facts are stored in the external memory. Consequently, selecting the relevant edits to accurately answering the questions becomes increasingly challenging for LLM-based QA.
@@ 4.3.1 Further Analysis
To evaluate the impact of KBQA on LLM-based QA within the GMeLLo framework, we conducted an analysis comparing the responses from LLMs to those from the KG. We consider the KG’s response as the final answer. Therefore, comparing to only using LLM-based QA, if the answer from LLMs is correct but the answer from the KG is incorrect, this leads to a decline in performance. Conversely, if the answer from LLMs is incorrect but the answer from the KG is correct, performance improves. If the KBQA provides no response, performance remains unchanged. As illustrated in Table 3, when there are discrepancies between KBQA and LLM-based QA responses, the likelihood of KBQA providing the correct answer increases as the parameter "k" increases.
@@
GPT-J-6B ${\color[rgb]{0,1,0}\boldsymbol{\uparrow}}{}$ 8.1 22.9 24.9 25.0 44.0 47.2 47.9 48.0
${\color[rgb]{1,0,0}\boldsymbol{\downarrow}}{}$ 12.5 2.4 1.2 0.7 0.7 0.4 0.3 0.3
@@
Vicuna-7B ${\color[rgb]{0,1,0}\boldsymbol{\uparrow}}{}$ 7.7 17.8 19.6 20.0 4.2 19.7 21.4 21.7
${\color[rgb]{1,0,0}\boldsymbol{\downarrow}}{}$ 21.8 3.9 2.0 1.2 7.2 4.2 4.0 3.9
@@ 4.4 Qualitative Analysis
Table 2 illustrates that Vicuna exhibits superior performance in directly handling the QA task, particularly when provided with the exact edited facts. Conversely, GPT-J excels in sentence analysis tasks, showcasing its high performance in the KBQA task.
@@ 4.4.1 Inferior Performance of GPT-J in QA
Table 2 shows that the performance of GPT-J and Vicuna in conducting QA tasks is comparable on the MQuAKE-CF dataset when k=1. However, GPT-J exhibits notably lower performance on the MQuAKE-T dataset. Further analysis revealed that GPT-J struggles in answering questions with only an edited fact pertaining to its intermediary information, such as:
{examplebox}
Sample from MQuAKE-CF
@@ 5 Conclusion
In this paper, we present GMeLLo, a method designed for multi-hop question answering in dynamic environments. In addition to leveraging LLMs for question answering, we also leverage the capabilities of LLMs to extract the triples from edited fact sentences to update a KG, and use the capabilities of LLMs to analyze question sentences and generate a relation chain, and finally get the formal query by filling in a formal query template. Finally, we combine KBQA and LLM-based QA to bolster the multi-hop question answering capability within a dynamic environment. This approach capitalizes on the strengths of both LLMs and KGs. By utilizing LLMs for analyzing question sentences and QA to ensure the coverage, and KBQA to provide accurate results, we achieve a synergy between these two methodologies.
@@
Acknowledgments
We thank all reviewers for providing valuable feedback. This work was partially supported by A*STAR CRF funding awarded to Cheston Tan. Bo Xiong is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB-1574 – 471687386.
@@ A.2 Hyperparmers Settings
To ensure reproducibility, we set the temperature to zero in all experiments. Table 5 shows that retrieving the top-6 edited facts from external memory provides the best average performance on the MQuAKE-CF dataset [^11] Tested only on the first question of each test instance, rather than all three for $k>1$. Consequently, we include top-6 edited facts in the prompt for subsequent experiments on this dataset when $k>1$. Similarly, for the MQuAKE-T dataset when $k>1$, we opted to incorporate the top-1 edited fact in the prompt.